A very loud source creates sound at the threshold of pain 1.5 m away. How far away, in km, will the sound just be audible to a human ear under ideal conditions?
Answer #1

130 = 10*log(I/Io)

So, I = 10^13*Io

Now, I/Io = 10^13

So, (r/1.5)^2 = 10^13

So, r = sqrt(10^13) *1.5 = 4.743*10^6 m

So, r = 4743 km